OpenStudy (anonymous):

help me please 10th grade math

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 2 points) Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions. Question 2 (Multiple Choice Worth 1 points) Stable High School has a total of 112 boys and girls who play sports. If the number of boys, b is 16 more than twice the number of girls, g, how many boys play sports at this high school? Which system of equations would help you solve this problem? b + g = 112 b = 2g + 16 b + g = 112 g = 2b + 16 b − g = 112 b = 2g − 16 None of the systems above would solve this problem. Question 3 (Multiple Choice Worth 2 points) Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1) Question 4 (Essay Worth 5 points) Part 1: If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system? Part 2: Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (cwtan):

you jz want to know how to get to the answer?

OpenStudy (anonymous):

i would like help understand these questions with answers please :)

OpenStudy (cwtan):

I explain question by question 1. from 1st eq we know that x=(1+y)/2 sub into eq 2 we get 2-2y+2y=-2 we get 2=-2 It is impossible so there is no solution for x.

OpenStudy (anonymous):

ok

OpenStudy (cwtan):

Question 2 Stable High School has a total of 112 boys and girls who play sports. We get B+G=112 the number of boys, b is 16 more than twice the number of girls, g, We get another eq B+16=2G or B=2G+16 Obviously there is no equation that can solve that question

OpenStudy (cwtan):

oops typing error

OpenStudy (cwtan):

B=2G-16 not +

OpenStudy (anonymous):

thanks for the help so far trying to finish quickly and study at the same time

OpenStudy (cwtan):

so did you still need me to explain the question left?

OpenStudy (cwtan):

Question 3 using 1st equation minus 2nd equation we get -2y-3y=14-9 solve for y then using the value y sub into any equation u will get value of x Note: the coordinate form is (x,y)

OpenStudy (anonymous):

yes i do

OpenStudy (cwtan):

Question 4 Part 1: The system of the of equation has it's solution. Part 2 Add the 2 equation together you will get 4y=20. similar as question 3 Get y and sub into any equation then you will found the value of x.

OpenStudy (anonymous):

how do i do this? ''Get y and sub into any equation then you will found the value of x''

OpenStudy (cwtan):

For Question 4 Part 2 Get y 4y=20__y=5 sub into y+2x=9 5+2x=9 2x=4 x=2

OpenStudy (anonymous):

what is the __ sorry for all the questions i just want to make sure i am able to do the exam

OpenStudy (cwtan):

sorry that is typing error it should be in next line

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks you were a huge help